How many liters of sulfur dioxide will be produced by the reaction of 1.1 mol of )2 at STP?


1
Expert's answer
2022-03-25T14:42:03-0400

How many liters of sulfur dioxide will be produced by the reaction of 1.1 mol of O2 at STP?

S + O2 = SO2

The mole ratio of SO2 to O is 1 : 1

So, moles SO2 = moles O2 = 1.1 mol

Volume = 1.1 * 22.4 = 24.64

(where: 22.4 L/mol – volume of 1 mole of any gas at STP)